OpenStudy (anonymous):

5(w + 2) – 11 = 2w + (9 – 7w)

OpenStudy (pounce0129):

5(w + 2) – 11 = 2w + (9 – 7w) 5w+10-11=2w+9-7w 5w-1=-5w+9 10w-1=9 10w=10 w=1